rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#1709 — FS#5690 — Switching filers - 2nd part

Attached to Project— VPS
Modernization
Storage
CLOSED
100%
Following the filer switching tasks performed on VPS, we are facing some particular cases.
Actually 39 VPS are causing problems:

vps10637 vps10666 vps13203 vps10319 vps10431
vps10498 vps17631 vps17816 vps17930 vps18601
vps10609 vps10971 vps11673 vps11755 vps10210
vps11092 vps11678 vps10420 vps10459 vps10804
vps10282 vps10912 vps11328 vps11580 vps10291
vps11749 vps11789 vps11923 vps11971 vps12031
vps17559 vps18680 vps10322 vps17625 vps11148
vps11156 vps11197 vps17887 vps11145

It is impossible to move quickly these servers since they dispose of a snapshot done by customers through a paid option.
In order to ensure the best service, we are going to contact personally the owners of these VPS to know whether it is possible to to delete their snapshots so we can switch without disrupting the service.
If ever the snapshot had an importance or in case we couldn't be able to contact the customer, we will be obliged to interrupt the service for some time to switch the disks (VPS switched off)

The deadline to contact these customers is fixed until Thursday the 25th August 2011.
After this date, VPS will be switched off on Friday the 26th August from 01h00 in the morning the time to move properly to the new filers.
The integrity of data is of certainly guaranteed. However, the bigger VPS disk is, longer the switching takes.


Date:  Friday, 26 August 2011, 02:38AM
Reason for closing:  Done
Comment by OVH - Wednesday, 17 August 2011, 16:20PM

We have succeeded in moving on the spot 19 VPS without any problem:

vps10319 vps10431 vps10498 vps17631 vps17816
vps17930 vps11092 vps11673 vps11145 vps11678
vps12031 vps11971 vps11789 vps11580 vps10459
vps10912 vps17625 vps11156 vps13203


However, we have a few problems with the host hosting the following vps:
vps18680 vps10322 vps11148 vps11197 vps17887

Le host ESXi refuses to be incorporated at the vCenter, and does not allow us to move the discs on the spot.
In order not to provoke an interruption of service, we prefer to leave this host on this status. These 5 VPS will obligatory be part of the forced updating to be carried out next week.

For the remaining les vps, we are still contacting their owners.


Comment by OVH - Wednesday, 17 August 2011, 16:21PM

7 more VPS were successfully migrated:

vps10637 vps10609 vps11755 vps10804 vps11749
vps10282 vps10291


Comment by OVH - Friday, 19 August 2011, 17:29PM

2 additional VPS were migrated : vps10210 vps10666


Comment by OVH - Friday, 26 August 2011, 00:14AM

The deadline for removal of snapshots has came now. The remaining VPS will then be migrated cold tonight from 1:00.

The VPS below will be impacted:

vps18601 vps10971 vps10420 vps11328 vps11923
vps17559 vps18680 vps10322 vps11148 vps11197
vps17887

The intervention will be performed as shown below :
- VPS will be shut down with the software way in the beginning(by OS) as long as we can.
Otherwise, they will be simply shut down on 'hardware'
- Then,the disks and their snapshots will be migrated on the new filers
- Eventually,VPS will be started.

In accordance with the size of the disk the intervention can last up to one hour per VPS. However, as there is especially 10G drives, we think it will be done before 3:00.
More informations will be available during the intervention.


Comment by OVH - Friday, 26 August 2011, 01:17AM

Updating started.


Comment by OVH - Friday, 26 August 2011, 02:38AM

The following VPS have been successfully switched:

vps18601 vps10971 vps11148 vps18680 vps10322
vps11197

Update is in process for the others.